OpenStudy (anonymous):

Find the eccentricity of the graph of ((x-2)^2/121) + ((y+6)^2/64)=1 and of the graph of ((x-2)^2/121)-((y+6)^2/64)=1 . Using complete sentences compare the eccentricity of the graph of an ellipse with the graph of a hyperbola.

OpenStudy (anonymous):

?? help??

OpenStudy (anonymous):

e=c/a for an ellipse c is the length of the semi major axis and a is the major axis since a is greater than b(the minor axis) and c a2=b2+c2 for hyperbola the same formula holds e=c/a but this time c is the largest thus c2=a2+b2 to make things easier try to collect variables together and see if it's an equation of an ellipse or hyperbola The first equation (x-2)^2/121) + ((y+6)^2/64)=1 is already simplified and is written in the form ((x-h)^2/a^2)+((y-k)^2/b^2)=1\ now from the given equation a^2=121 and b^2=64 thus the ellipse is an x (horizontal ellipse) to find c use the above formula I gave you a^2=b^2+c^2 c=sqrt(a^2-b^2) c=sqrt(121-68) c=sqrt(53) and because a^2=121 a=sqrt(121) e=c/a=(sqrt(53))/(sqrt(121)) use similar steps to find the ecentricity for the hyperbola.but remeber that c(the transverse axis is the largest so instead of the formula a^2=b^2+c^2 you should use c^2=a^2+^2) Let me know if you have any doubts.